# Binary Black Hole Waveforms from High-Resolution GR-Athena++ Simulations

### v1

@@ Abstract
The detection and subsequent inference of binary black hole signals rely heavily on the accuracy of the waveform model employed. In the highly non-linear, dynamic, and strong-field regime near merger, these waveforms can only be accurately modeled through numerical relativity simulations. Considering the precision requirements of next-generation gravitational wave observatories, we present in this paper high-resolution simulations of four non-spinning quasi-circular binary black hole systems with mass ratios of 1, 2, 3, and 4, conducted using the GR-Athena++ code. We extract waveforms from these simulations using both finite radius and Cauchy characteristic extraction methods. Additionally, we provide a comprehensive error analysis to evaluate the accuracy and convergence of the waveforms. This dataset encompasses gravitational waves of the precision (self-mismatch) demanded by upcoming gravitational detectors such as LISA, Cosmic Explorer, and Einstein Telescope. The waveforms are publicly available on ScholarSphere [1, 2, 3, 4], and represent the first set of waveforms of the new GR-Athena++ catalog.
@@ I Introduction
The rest of this paper is structured as follows. In Sec. II, we provide a detailed overview of the numerical methods, including grid configurations, GW extraction techniques, and error estimation models. In Sec. III we present our simulations, showcasing the waveforms for different modes along with the trajectories of the BBH systems for each mass ratio. Sec. III is dedicated to evaluating the quality of the waveforms through a series of convergence studies and cross-validation against waveforms from the SXS catalog. Specifically, we analyse the convergence of Cauchy characteristic extraction (CCE) strains and estimate their errors contributed from different parameters such as the resolution and world tube extraction radius. Comparisons against SXS are performed via time-domain alignment as well as frequency-domain mismatch calculations, considering both single-mode and multi-mode waveforms. Finally, Sec. IV concludes our work with an overview of our results and reflections on future research directions. Additional material for further examination is provided in the appendices.
@@ II Numerical simulations
Figure 4: Sum of errors in CCE strains (computed from $\psi_{4}^{\ell m}$). $err_{\mathrm{PIT}}(N^{-p})$ denotes the error for different resolutions of the PITTNull code, and $p$ is the convergence order $err_{\mathrm{PIT}}(R^{-p})$ is the error for different world tube radii of the PITTNull code; $err_{\mathrm{GRA}}(N^{-p})$ shows error pertinent to GR-Athena++’s resolutions. The total error $err$ for $\phi$ and amplitude relative difference $\Delta A/A$ is computed by $err=(err^{2}_{\mathrm{PIT}}(N^{-p})+err^{2}_{\mathrm{PIT}}(R^{-p})+err^{2}_{\mathrm{GRA}}(N^{-p}))^{1/2}$.
@@ II.2 Grid setup
The resolution of the root (or base) grid is determined by specifying the number of grid points in each direction. We divide the root grid into MeshBlocks, each containing a predetermined number of grid points that must be a divisor of the total points in the root grid. These MeshBlocks are recursively subdivided to create the full adaptive mesh refinement octree structure grid, starting from the root grid.
@@
In our simulations, the MeshBlock’s size, in all directions, is set to $16$. The number of refinement levels, set between 12 and 14, is chosen based on the BBH’s mass ratio, to ensure a minimum resolution of $m_{2}/\delta{x}>25$ at the lowest resolution of the root grid, where $m_{2}$ is the mass of lighter or secondary BH and $\delta{x}$ is the grid spacing at the puncture.
@@ II.4 Gravitational waves
To compute waveforms at null infinity $\mathcal{I}_{+}$, we use either an approximate method based on post-Newtonian theory (finite radius extraction; FRE), or the CCE method. The latter directly solves Einstein’s equations to $\mathcal{I}_{+}$, and therefore represents our preferred technique for computing gravitational waves. We describe it in detail in this section. Additional information regarding the finite radius extraction (FRE) method can be found in App. (A), while App. (D) collects comparisons of data obtained with the two methods to evaluate their systematic uncertainties.
@@
App. (E) shows, as an example, the $C_{222}$ coefficients for $\gamma_{xx}$. The presence of noise in the $C_{klm}$ coefficients at the beginning of the run and even at a later times is one of the challenges to deal with when computing CCE strains. We have tested different filtering methods to address this issue, as detailed in App. (E), but ultimately chose not to filter $C_{klm}$ when computing the presented waveforms.
@@
where the dots indicate time derivatives. In both scenarios, the initial step involves normalizing the quantities by multiplying them by a factor of $2\times(-1)^{l+m}$. To recover the strain from either, we use the fixed frequency integration (FFI) method [87]. The cutoff frequency employed for this process is given in App. (B).
@@ II.5 Uncertainty quantification
We estimate the finite-resolution and finite-extraction radius errors in our NR simulations by performing a series of convergence studies on the strain data obtained using CCE, as discussed in Sec. III.2, and FRE, as detailed in App. (C). In both cases, we model the error in any numerically computed quantity $w$ as a power of $\delta$ [88, 89]
@@ III.1 Simulations
The trajectories for each puncture, $\vec{x}_{p}(t)$, are found by solving the ordinary differential equation $\dot{\vec{x}}_{p}(t)=-\vec{\beta}_{p}(t)$, in which, $\vec{\beta}_{p}(t)$ denotes the shift vector at the puncture [21, 7]. We estimate the eccentricity of our simulations shown in Fig. 1 using the methods discussed in [91]. This approach requires us to identify parameters, including eccentricity, to fit the coordinate distance between punctures, derived from the trajectories $\vec{x}_{p}(t)$, as a function of evolution time. We use the initial orbital angular velocity and the first few orbits of the simulation to find a fit for the coordinate distance versus time. We choose the coordinate distance and not its time derivative, as done in [92], because the time derivative of the coordinate distance can introduce noise. In particular, we use $\sim 4$ orbits and calculate the initial orbital velocity using the initial momenta and distance between the puncture, see App. (B). Specifically, BBH runs with $q=$ 2, 3, 4 have an eccentricity of $\sim 10^{-3}$, and the $q=1$ case has an eccentricity of $\sim 6\times 10^{-4}$.
Fig. (2) shows the $(\ell=2,m=2)$ and $(\ell=4,m=4)$ modes of the CCE strains for our simulations. To ensure that the merger happens at $t=0$, we subtract the merger time $t_{\mathrm{mrg}}$, defined as the peak amplitude time of the $(2,2)$ mode, in the time axis. At $t_{\mathrm{mrg}}$, a highly excited BH is formed, which then settles down to its stationary Kerr state by emitting GWs. In Fig. (2), we see each system exhibits a clear inspiral phase, followed by a merger and a ringdown. The subdominant $(4,4)$ mode also captures these key features, albeit with a smaller amplitude.
@@ III.2 CCE error analysis
After computing the CCE strains from $\psi_{4}^{\ell m}$, using FFI (App. (B)), we want to determine the convergence order, and hence finding an estimate of the error budget in the waveforms.
@@
We first find the convergence order with respect to the PITTNull resolution. Accordingly, we use Eq. (5) with $\delta=N^{-1}$, where $N$ refers to the $N_{r}$ resolution of PITTNull runs. Fig. (3) depicts the convergence of CCE strain’s phase computed from $\psi_{4}^{\ell m}$ with a world tube radius of $50$  [^1] Among different world tube extraction radii, only radius = $50$ exhibits convergence behavior with respect to PITTNull resolutions.. We observe a 2nd order convergence with respect to the PITTNull resolution. Similary, we conduct a similar convergence study for CCE strain’s amplitude and find a 2nd order convergence in all runs except $q=1$ that exhibits a 1st order convergence.
For the convergence order with respect to the PITTNull extraction world tube, we use the highest resolution of GR-Athena++ and PITTNull, and use Eq. (5) with $\delta=R^{-1}$, in which $R$ refers to the world tube radius extraction of $20$, $50$, and $100$. The result of this convergence study shows that the CCE strain’s phase demonstrates a consistent 2nd order convergence for all mass ratios. However, the amplitude converges in different orders for different mass ratios. In particular, the amplitude in $q=1$ shows no convergence, $q=2$ converges with a 2nd order rate, and $q=3$ and 4 converge with a 3rd order rate.
@@
We summarize the results of the convergence study as well as the error analyses, for both amplitude and phase of the (2,2) mode, in Fig. (4). We show the error from different sources: the PITTNull resolution $err_{\mathrm{PIT}}(N^{-p})$, the extraction radius $err_{\mathrm{PIT}}(R^{-p})$, and the GR-Athena++ resolution $err_{\mathrm{GRA}}(N^{-p})$, where $p$ denotes the convergence order. Additionally, we represent the total error with a solid blue line. This total error is calculated as the quadrature sum of the individual errors.
@@
In Fig. (4), we observe that the PITTNull resolution error for the phase, denoted as $err_{\mathrm{PIT}}(N^{-p})$, is the smallest error among various sources of error for all mass ratios except $q=1$. For amplitude, $err_{\mathrm{PIT}}(N^{-p})$ is smallest only for mass ratios $q=3$ and $q=4$. The phase error shows a slight increase as we approach the merger, ranging from $(10^{-5},\sim 10^{-3})$ for different mass ratios. In contrast, the error in amplitude remains relatively constant for $q=$ 2, 3, and 4 cases, and it is slightly increasing only for $q=1$. The range of this error is in $(10^{-4},\sim 5\times 10^{-2})$.
@@
The error of the extraction radius is denoted by $err_{\mathrm{PIT}}(R^{-p})$ in Fig. (4). We observe that the largest source of error is due to the extraction radius for both phase and amplitude. To mitigate this error, one can use a larger extraction radius. For example, by using a radius of $100$, the error term reduces by a factor of 4, cf. Eq. (7). However, only the strains with the extraction radius of 50 show a consistent convergence order. Moreover, as we discuss in Sec. (III.3), the strains with the extraction radius of $R=50$ exhibit much smaller mismatch than $R=100$. Consequently, we select to use strains with world tube execration $R=50$ for our analysis.
@@
In Fig. (4), we see that the error in the phase and amplitude associated to the GR-Athena++ resolution, $err_{\mathrm{GRA}}(N^{-p})$, is relatively larger during the inspiral for high mass ratio runs compared to lower mass ratio runs with $q=$ 1 and 2. Moreover, this error grows as it gets close to the merger, and it is within the range of $(10^{-5},10^{-2})$ for both phase and amplitude.
@@
For $q=$ 2, 3, and 4, we observe in Fig. (4) that during the inspiral, and after junk radiation, the total error for the dephasing is less than 0.01 radians. Moreover, the total error for the amplitude relative difference is below $\sim 1\%$. As we get close to the merger, the total error for these cases increases but still remains below 0.05 radians for the phase and below $\sim 5\%$ for the amplitude relative difference. For $q=1$, we see a larger total error, in its phase and amplitude, with respect to the higher mass ratio runs.
@@ III.3 Self-mismatch
Our goal is to assess the mismatch of the highest resolution run in table (1) against the ‘exact’ benchmark, i.e., an infinite resolution run, for the next-generation detectors such as LISA [95], the ET [62], and CE [59]. To this aim, we first model the mismatch between lower resolutions with respect to the highest available resolution in table (1) using
@@
Fig. (5) shows the mismatch of the $(2,2)$ mode strains obtained using the CCE method for two different world tube extraction radii of 50, the left panel, and 100, the right panel. We employ the LISA noise curve [97] for this comparison and consider a total mass of $10^{5}M_{\odot}$. The discrete data points represent the mismatch between the highest and the other resolutions. Additionally, the corresponding fit, described by Eq. (13) with $\delta x_{\mathrm{ref}}\to 0$, is shown as a dashed line, with different colors indicating each mass ratio. An upside-down triangle on the dashed line indicates the mismatch between the highest resolution and the exact solution, i.e., evaluating Eq. (13) with the fitted values of $c$ and $p$ and setting $\delta x_{\mathrm{r}ef}=0$.
@@
We see in the left panel of Fig. (5) that the mismatch between the highest resolution run and the second highest resolution run at $R=50$ is $\sim 5\times 10^{-9}$ for all mass ratios. Additionally, we observe that for mass ratios $q=1$, $q=2$, and $q=3$, the mismatch of our highest resolution strain against the “exact” strain is $\sim 10^{-12}$, while the $q=4$ run exhibits a mismatch $\sim 10^{-11}$. The latter mismatches are orders of magnitude more accurate than the $\sim 10^{-7}$ threshold required by LISA.
In the right panel of Fig. (5), we present the mismatch between the waveforms extracted at $R=100$. The mismatch between the highest resolution strain and the second highest resolution strain is $\sim 3\times 10^{-5}$ for $q=$ 1, 2, and 3 runs, and $\sim 10^{-4}$ for $q=4$. The mismatch of the highest resolution strain against the “exact” strain, namely, when $\delta x_{\mathrm{r}ef}\to 0$ in Eq. (13), is $\sim 10^{-5}$ for mass ratios of 1, 2, and 3, while it is $\sim 10^{-4}$ for the $q=4$ run. These mismatches are significantly larger than those observed in the left panel and do not meet the mismatch requirement set by LISA.

### v3

@@ Abstract
The detection and subsequent inference of binary black hole signals rely heavily on the accuracy of the waveform model employed. In the highly non-linear, dynamic, and strong-field regime near merger, these waveforms can only be accurately modeled through numerical relativity simulations. Considering the precision requirements of next-generation gravitational wave observatories, we present in this paper high-resolution simulations of four non-spinning quasi-circular binary black hole systems with mass ratios of 1, 2, 3, and 4, conducted using the GR-Athena++ code. We extract waveforms from these simulations using both finite radius and Cauchy characteristic extraction (CCE) methods. Additionally, we provide a comprehensive error analysis to evaluate the accuracy and convergence of the waveforms. Our self-mismatch study shows that the (2, 2) mode of the CCE strains, for the world tube extraction radius of $R=50$, reaches the level of $\sim 10^{-12}$ mismatch for mass ratios of 1, 2, 3, and $\sim 10^{-11}$ mismatch for the mass ratio of 4. However, when larger extraction radii are considered or when more modes are included the mismatches increase. These results highlight both the promise and limitations of current simulations in achieving the precision required for upcoming detectors such as LISA, Cosmic Explorer, and Einstein Telescope. The waveforms are publicly available on ScholarSphere [1, 2, 3, 4], and represent the first set of waveforms of the new GR-Athena++ catalog.
@@ I Introduction
The rest of this paper is structured as follows. In Sec. II, we provide a detailed overview of the numerical methods, including grid configurations, GW extraction techniques, and error estimation models. In Sec. III we present our simulations, showcasing the waveforms for different modes along with the trajectories of the BBH systems for each mass ratio. Sec. III is dedicated to evaluating the quality of the waveforms through a series of convergence studies and cross-validation against waveforms from the SXS catalog. Specifically, we analyse the convergence of CCE strains and estimate their errors contributed from different parameters such as the resolution and world tube extraction radius. Comparisons against SXS are performed via time-domain alignment as well as frequency-domain mismatch calculations, considering both single-mode and multi-mode waveforms. Finally, Sec. IV concludes our work with an overview of our results and reflections on future research directions. Additional material for further examination is provided in the appendices.
@@ II Numerical simulations
Figure 4: Convergence of CCE strains, (2,2) modes with respect to the PITTNull’s world tube extraction radius. Here, we have $\delta_{i}^{-1}\in[20,50,100]$ for $i=1,2,3$. The strains are computed from $\psi_{4}^{\ell m}$. The boundary condition for PITTNull is obtained by the highest resolution from the table (1) and it is run using the highest resolution of the PITTNull code, see the text.
Figure 5: Sum of errors for the $(2,2)$ mode of CCE strains (computed from $\psi_{4}^{\ell m}$). $err_{\mathrm{PIT}}(N^{-p})$ denotes the magnitude of the leading order error term, Eq. (7), at the highest resolution of PITTNull run, and $p$ is the convergence order. Similarly, $err_{\mathrm{PIT}}(R^{-p})$ denotes the error due to world tube extraction radius at $R=50$; $err_{\mathrm{GRA}}(N^{-p})$ marks the error pertinent to GR-Athena++’s resolutions, at the highest available resolution of GR-Athena++. The total error $err$ for $\phi$ and amplitude relative difference $\Delta A/A$ is computed by $err=(err^{2}_{\mathrm{PIT}}(N^{-p})+err^{2}_{\mathrm{PIT}}(R^{-p})+err^{2}_{\mathrm{GRA}}(N^{-p}))^{1/2}$.
@@ II.2 Grid setup
The resolution of the root (or base) grid is determined by specifying the number of grid points in each direction. We divide the root grid into MeshBlocks, each containing a predetermined number of grid points that must be a divisor of the total points in the root grid. Consequently, for a fixed grid extent, increasing the root grid resolution generates more MeshBlocks at the root grid thereby increasing the resolution of each MeshBlock. Then, these MeshBlocks are recursively subdivided to create the full adaptive mesh refinement octree structure grid, starting from the root grid.
@@
In our simulations, the MeshBlock’s size, in all directions, is set to $16$. The number of refinement levels, set between 12 and 14, is chosen based on the BBH’s mass ratio, to ensure a minimum resolution of $m_{2}/\delta{x}>25$ at the lowest resolution of the root grid, where $m_{2}$ is the mass of lighter or secondary BH and $\delta{x}$ is the grid spacing at the puncture. Additionally, the 2:1 refinement constraint further improves resolution in the gravitational wave (GW) extraction regions by increasing the refinement level.
@@ II.4 Gravitational waves
To compute waveforms at null infinity $\mathcal{I}_{+}$, we use either an approximate method based on post-Newtonian theory (finite radius extraction; FRE), or the CCE method. The latter directly solves Einstein’s equations to $\mathcal{I}_{+}$, and therefore represents our preferred technique for computing GWs. We describe it in detail in this section. Additional information regarding the finite radius extraction (FRE) method can be found in App. A, while App. D collects comparisons of data obtained with the two methods to evaluate their systematic uncertainties.
@@
App. E shows, as an example, the $C_{222}$ coefficients for $\gamma_{xx}$. The presence of noise in the $C_{klm}$ coefficients at the beginning of the run and even at a later times is one of the challenges to deal with when computing CCE strains. We have tested different filtering methods to address this issue, as detailed in App. E, but ultimately chose not to filter $C_{klm}$ when computing the presented waveforms.
@@
where the dots indicate time derivatives. In both scenarios, the initial step involves normalizing the quantities by multiplying them by a factor of $2\times(-1)^{l+m}$. To recover the strain from either, we use the fixed frequency integration (FFI) method [87]. The cutoff frequency employed for this process is given in App. B.
It is important to note that while our CCE strains include GW memory effects [88], the evaluation of these effects requires the complete past history of the spacetime. During the FFI process, we lose the integration constant that represents the history of the mode prior to the start of the NR simulation. Consequently, the FFI method does not account for this constant of integration related to the memory effect. Additionally, our choice of cutoff frequency is optimized for the oscillatory modes of GW strains. Readers may refer to [89], where GW strains are matched with post-Newtonian strains to incorporate GW memory effects. Furthermore, we do not apply any supertranslations to transform the strains into the appropriate BMS frame, as discussed in [90].
@@ II.5 Uncertainty quantification
We estimate the finite-resolution and finite-extraction radius errors in our NR simulations by performing a series of convergence studies on the strain data obtained using CCE, as discussed in Sec. III.2, and FRE, as detailed in App. C. In both cases, we model the error in any numerically computed quantity $w$ as a power of $\delta$ [91, 92]
@@ III.1 Simulations
The trajectories for each puncture, $\vec{x}_{p}(t)$, are found by solving the ordinary differential equation $\dot{\vec{x}}_{p}(t)=-\vec{\beta}_{p}(t)$, in which, $\vec{\beta}_{p}(t)$ denotes the shift vector at the puncture [21, 7]. The initial values of $\vec{x}_{p}(t=0)$ are used from the initial data parameters, and hence by solving this ordinary differential equation, we can find the trajectory of each BH during the evolution. This method is a standard method and also used in e.g., Einstein Toolkit [7]. Additionally, in AthenaK [94], we have experimented with tracking the minimum of the lapse on the grid, but it is less accurate than solving the the ordinary differential equation for $\vec{x}_{p}(t)$.
We estimate the eccentricity of our simulations shown in Fig. 1 using the methods discussed in [95]. This approach requires us to identify parameters, including eccentricity, to fit the coordinate distance between punctures, derived from the trajectories $\vec{x}_{p}(t)$, as a function of evolution time. We use the initial orbital angular velocity and the first few orbits of the simulation to find a fit for the coordinate distance versus time. We choose the coordinate distance and not its time derivative, as done in [96], because the time derivative of the coordinate distance can introduce noise. In particular, we use $\sim 4$ orbits and calculate the initial orbital velocity using the initial momenta and distance between the puncture, see App. B. Specifically, BBH runs with $q=$ 2, 3, 4 have an eccentricity of $\sim 10^{-3}$, and the $q=1$ case has an eccentricity of $\sim 6\times 10^{-4}$.
Fig. 2 shows the $(\ell=2,m=2)$ and $(\ell=4,m=4)$ modes of the CCE strains for our simulations. To ensure that the merger happens at $t=0$, we subtract the merger time $t_{\mathrm{mrg}}$, defined as the peak amplitude time of the $(2,2)$ mode, in the time axis. At $t_{\mathrm{mrg}}$, a highly excited BH is formed, which then settles down to its stationary Kerr state by emitting GWs. In Fig. 2, we see each system exhibits a clear inspiral phase, followed by a merger and a ringdown. The subdominant $(4,4)$ mode also captures these key features, albeit with a smaller amplitude.
@@ III.2 CCE error analysis
After computing the CCE strains from $\psi_{4}^{\ell m}$, using FFI (App. B), we want to determine the convergence order, and hence finding an estimate of the error budget in the waveforms.
@@
We first find the convergence order with respect to the PITTNull resolution. Accordingly, we use Eq. (5) with $\delta=N^{-1}$, where $N$ refers to the $N_{r}$ resolution of PITTNull runs. Fig. 3 depicts the convergence of CCE strain’s phase computed from $\psi_{4}^{\ell m}$ with a world tube radius of $50$  [^1] Among different world tube extraction radii, only radius = $50$ exhibits convergence behavior with respect to PITTNull resolutions.. We observe a 2nd order convergence with respect to the PITTNull resolution. Similary, we conduct a similar convergence study for CCE strain’s amplitude and find a 2nd order convergence in all runs except $q=1$ that exhibits a 1st order convergence.
For the convergence order with respect to the PITTNull extraction world tube, we use the highest resolution of GR-Athena++ and PITTNull, and use Eq. (5) with $\delta=R^{-1}$, in which $R$ refers to the world tube radius extraction of $20$, $50$, and $100$. The result of this convergence study, depicted in Fig. 4, shows that the CCE strain’s phase demonstrates a consistent $\sim$ 2nd order convergence for all mass ratios. However, the amplitude converges in different orders for different mass ratios. In particular, the amplitude in $q=1$ shows no convergence, $q=2$ converges with a 2nd order rate, and $q=3$ and 4 converge with a 3rd order rate.
@@
We summarize the results of the convergence study as well as the error analyses, for both amplitude and phase of the (2,2) mode, in Fig. 5. Specifically, using Eq. (7), we estimate the magnitude of the leading order error term, expanded in PITTNull’s resolutions, at the highest resolution of PITTNull run; we denote this error as $err_{\mathrm{PIT}}(N^{-p})$, where $p$ shows the convergence order. Similarly, we denote the error due to radius extraction at $R=50$ as $err_{\mathrm{PIT}}(R^{-p})$, and, the error due to GR-Athena++’s resolutions, at the highest resolution run of GR-Athena++, as $err_{\mathrm{GRA}}(N^{-p})$. Additionally, we represent the total error with a solid blue line. This total error is calculated as the quadrature sum of the individual errors.
@@
In Fig. 5, we observe that the PITTNull resolution error for the phase, denoted as $err_{\mathrm{PIT}}(N^{-p})$, is the smallest error among various sources of error for all mass ratios except $q=1$. For amplitude, $err_{\mathrm{PIT}}(N^{-p})$ is smallest only for mass ratios $q=3$ and $q=4$. The phase error shows a slight increase as we approach the merger, ranging from $(10^{-5},\sim 10^{-3})$ for different mass ratios. In contrast, the error in amplitude remains relatively constant for $q=$ 2, 3, and 4 cases, and it is slightly increasing only for $q=1$. The range of this error is in $(10^{-4},\sim 5\times 10^{-2})$. The observed convergence, with respect to the PITTNull’s resolution, suggests $err_{\mathrm{PIT}}(N^{-p})$ decreases as the resolution increases. For instance, Ref. [97] finds a maximum error of $\sim 10^{-6}$ in a bouncing BH system when the resolution of $900\times 900\times 900$ was used in PITTNull.
@@
The error of the extraction radius is denoted by $err_{\mathrm{PIT}}(R^{-p})$ in Fig. 5. We observe that the largest source of error is due to the extraction radius for both phase and amplitude. To mitigate this error, one can use a larger extraction radius. For example, by using a radius of $100$, the error term reduces by a factor of 4, cf. Eq. (7). However, only the strains with the extraction radius of 50 show a consistent convergence order. Moreover, as we discuss in Sec. (III.3), the strains with the extraction radius of $R=50$ exhibit much smaller mismatch than $R=100$. Consequently, we select to use strains with world tube execration $R=50$ for our analysis. It is worth noting that while theoretically the CCE strains are solutions to the full Einstein equations, and thus should not depend on the extraction radius, practical considerations complicate this expectation. Specifically, in the PITTNull code, there is an assumption [16] regarding the geometry of the initial null hypersurface, which is presumed to be close to a Schwarzschild geometry. This assumption is necessary for setting the initial characteristic data, making it less accurate for smaller extraction radii. Conversely, larger radii can introduce challenges due to noise present in our data, as discussed in App. E, as well as numerical errors arising from small corrections to a flat geometry which may not be adequately captured by our Cartesian grid.
@@
In Fig. 5, we see that the error in the phase and amplitude associated to the GR-Athena++ resolution, $err_{\mathrm{GRA}}(N^{-p})$, is relatively larger during the inspiral for high mass ratio runs compared to lower mass ratio runs with $q=$ 1 and 2. Moreover, this error grows as it gets close to the merger, and it is within the range of $(10^{-5},10^{-2})$ for both phase and amplitude.
@@
For $q=$ 2, 3, and 4, we observe in Fig. 5 that during the inspiral, and after junk radiation, the total error for the dephasing is less than 0.01 radians. Moreover, the total error for the amplitude relative difference is below $\sim 1\%$. As we get close to the merger, the total error for these cases increases but still remains below 0.05 radians for the phase and below $\sim 5\%$ for the amplitude relative difference. For $q=1$, we see a larger total error, in its phase and amplitude, with respect to the higher mass ratio runs.
@@ III.3 Self-mismatch
Our goal is to assess the mismatch of the highest resolution run in table (1) against the ‘exact’ benchmark, i.e., an infinite resolution run, for the next-generation detectors such as LISA [100], the ET [62], and CE [59]. Whenever we evaluate the mismatches for these detectors, we employ the frequency intervals given by the starting frequency of the GR-Athena++ waveform $f_{\ell m}^{0}=m/2\omega_{22}/(2\pi)$ (multiplied by 1.25, to avoid artifacts due to FFTs) and $1024$ or $1$ Hz, respectively for CE and LISA. Additionally, when higher modes are considered, we fix the inclination angle to $\iota=\pi/3$. To this aim, we first model the mismatch between lower resolutions with respect to the highest available resolution in table (1) using
@@
Fig. 6 shows the mismatch of the $(2,2)$ mode strains obtained using the CCE method for two different world tube extraction radii of 50, the left panel, and 100, the right panel. We employ the LISA noise curve [102] for this comparison and consider a total mass of $10^{5}M_{\odot}$. The discrete data points represent the mismatch between the highest and the other resolutions. Additionally, the corresponding fit, described by Eq. (13) with $\delta x_{\mathrm{ref}}\to 0$, is shown as a dashed line, with different colors indicating each mass ratio. An upside-down triangle on the dashed line indicates the mismatch between the highest resolution and the exact solution, i.e., evaluating Eq. (13) with the fitted values of $c$ and $p$ and setting $\delta x_{\mathrm{r}ef}=0$.
@@
We see in the left panel of Fig. 6 that the mismatch between the highest resolution run and the second highest resolution run at $R=50$ is $\sim 5\times 10^{-9}$ for all mass ratios. Additionally, we observe that for mass ratios $q=1$, $q=2$, and $q=3$, the mismatch of our highest resolution strain against the “exact” strain is $\sim 10^{-12}$, while the $q=4$ run exhibits a mismatch $\sim 10^{-11}$. The latter mismatches are orders of magnitude more accurate than the $\sim 10^{-7}$ threshold required by LISA.
In the right panel of Fig. 6, we present the mismatch between the waveforms extracted at $R=100$. The mismatch between the highest resolution strain and the second highest resolution strain is $\sim 3\times 10^{-5}$ for $q=$ 1, 2, and 3 runs, and $\sim 10^{-4}$ for $q=4$. The mismatch of the highest resolution strain against the “exact” strain, namely, when $\delta x_{\mathrm{r}ef}\to 0$ in Eq. (13), is $\sim 10^{-5}$ for mass ratios of 1, 2, and 3, while it is $\sim 10^{-4}$ for the $q=4$ run. These mismatches are significantly larger than those observed in the left panel and do not meet the mismatch requirement set by LISA.